40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 190 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  7517   Wed Oct 10 08:36:47 2012 SteveUpdateCOCspecial mirror mounts holder

Quote:

After looking at the in-vacuum layout I think we should make two changes during the next vent:

1) Reduce the number of mirrors between the FI and its camera. We install a large silvered mirror in the vacuum flange which holds the Faraday cam (in the inside of the viewport). That points directly at the input to the Faraday. We get to remove all of the steering mirror junk on the IO stack.

2) Take the Faraday output (IFO REFL) out onto the little table holding the BS and PRM Oplevs. We then relocate all 4 of the REFL RFPDs as well as the REFL OSA and the REFL camera onto this table. This will reduce the path length from the FI REFL port to the diodes and reduce the beam clutter on the AS table.

1)  Mirror mount holder for "large silvered mirror" inside of the 8" OD tube vacuum envelope.

Attachment 1: 10101201.PDF
10101201.PDF
  7516   Wed Oct 10 02:20:34 2012 ranaUpdateSUSOptical Lever QPD mods

 Since we upgraded the CDS system, I guess our ADC ranges have gone up but we never did anything to the OLs to match the ADC ranges. From Liz's daily summary page of the OL, I see this:

C1-CORE_OPTICS_SUM_DATA_ALL_TIME_0-1033801216-86400.png

So we need a factor of 5-10 increase in the electronics gain (why isn't the BS SUM on there?). This might be accomplished in the head, but for the ones with whitening boards, might be better to do there.

(** add to Jamie's list of long term tasks **)

  7515   Wed Oct 10 02:15:14 2012 ranaUpdateLSC11 MHz reconnected to EOM

 Absolutely hokey. What are our requirements for this RFPD? What are the power levels and SNR that we want (I seem to remember that its for 22 as well as 110 MHz)? Perhaps we can test an aLIGO one if Rich has one sitting around, or if the aLIGO idea is to use a broadband PD I guess we can just keep using what we have.

  7514   Wed Oct 10 00:18:58 2012 JenneUpdateLockingREFL camera aligned

I moved some of the REFL optics on the AS table by a teeny bit to accomodate the new place that the REFL beam exits the chamber (none of this was done while we were at air....we were only dealing with the AS beam at the time, and were happy that REFL came out of the vacuum).

The REFL beam is now on the REFL camera (with PRMI aligned), and the beam is going toward the 4 REFL RF PDs, but it's not aligned to any of them.

I have some questions as to mystery optics on in the REFL path.  There is a 90% BS, and I don't know where the 10% reflection goes....is it going to beat against the AUX Stochino laser?

I have to go, and I didn't fix the videocapture script today, so pix tomorrow, I promise.

  7513   Tue Oct 9 23:12:56 2012 JenneUpdateLSC11MHz reconnected to EOM

Riju hasn't been in the lab in a long time to do any measurements, so I put the signals back to how they should be. 

I turned off / confirmed off the things which were sending signal to the EOM:  the network analyzer, the RF generator box, and the Marconi which supplies the 11MHz. 

I removed the cavity scanning cable, and terminated it, and put the regular 11MHz cable back on the splitter.

I then turned on the RF gen box and the Marconi.  The Marconi had been off, so we were not getting any 11MHz or 55MHz out of the RF gen. box.  This is why I couldn't lock any cavities last night (duh). 

On to locking!

----------------- In other news,

While swapping out the EOM cable, I noticed that the DC power supply sitting under the POX table was supplying a weird value, 17 point something volts.  I checked on the table to remind myself why that power supply is there...it's powering an RF amplifier right after the commercial PD that is acting as POP22.  The amplifier wants +15 and GND, so I reset the power supply to 15V.  We should add this to the list of things to fix, because it's dumb.  Either we need to put in the real POP22 (long term goal), or we need to get this guy some rack power, and do the same for any amplifiers for the Beat setup.  It's a little hoakey to have power supplies littering the lab.

  7512   Tue Oct 9 17:33:37 2012 jamieUpdateSUSdiagonalization

Quote:

I went inside to align the beam on WFS and noticed that oscillations in yaw are ~10 times stronger then in pitch. I've plot rms of pitch and yaw measured by LID sensors and saw that MC3 yaw rms motion is a few times larger then pitch.

What are "LID" sensors?  Do you mean the OSEM shadow sensors?  I'm pretty sure that's what you meant, but I'm curious what "LID" means.

 

  7511   Tue Oct 9 17:16:14 2012 DenUpdateSUSdiagonalization

I went inside to align the beam on WFS and noticed that oscillations in yaw are ~10 times stronger then in pitch. I've plot rms of pitch and yaw measured by LID sensors and saw that MC3 yaw rms motion is a few times larger then pitch.

Also MC1 input diag matrix does not diagonalize signals for pitch and yaw. In the spectrums of these signals all 4 resonance are equally seen during the free swinging. I think we should rediagonalize MC1.

Another thing is that if MC1 and MC3 are on the same stack,  pitch and yaw spectrums of these mirrors should be comparable. But MC1 signal is ~2-3 times larger then of MC3. I think we should correct calibration.

Attachment 1: mc123_rms.pdf
mc123_rms.pdf
Attachment 2: mc1_diag.pdf
mc1_diag.pdf
Attachment 3: mc123_rms.pdf
mc123_rms.pdf
  7510   Tue Oct 9 09:29:10 2012 SteveUpdateIOOthis is where we are

Quote:

Quote:

 The new cold cathode gauge CC1 is in place. We were at atmosphere for 28 days ......more later

 

 cc1 = 2.3e-5 Torr at day 6 vacuum normal

 

Attachment 1: 15days.png
15days.png
  7509   Tue Oct 9 00:25:33 2012 JenneUpdateIOOPZTs - hacky solution in place!!

[Evan, Jenne]

We applied some volts across both the pitch and yaw pin sets of the ribbon cable that goes to PZT2.  We ended up with ~40V yaw and ~14.5V pitch.  That was the nice happy center of the clipping that we can see on the AS camera.  Once we found the center of the PZT clipping range with the ITMY beam, we recentered the AS camera (actually, this took a few iterations, but now it's good). 

We then aligned MICH, but aren't able to get it to lock.  Before falling asleep, we have decided to align the PRM and SRM, so right now we have a flashing DRMI.  Both the SRMI and PRMI look a little funny the closer you get to 'good' alignment, so I'll investigate a little more tomorrow, and include pictures.  (The video capture script has barfed again, and I'm not in the mood to deal with it today.)

  7508   Mon Oct 8 23:58:57 2012 JenneUpdateSAFETYControl room emergency lights came on

[Evan, Jenne]

We were sitting trying to lock MICH (hooooorraaaayy!!!), and the emergency lights above the control room door came on, and then ~1 minute later turned off.  Steve, can you see what's up?

  7507   Mon Oct 8 22:07:46 2012 ranaUpdateSUSDoors on, ready to pump

Quote:

All oplevs need a little realignment, especially ETMY, which had it's lens removed (Rana has a Wall of Shame photo of this, which is why it was removed by him).  Steve will look into this tomorrow, after he starts pumping.

 The shame:

20120921_200637.jpg

There is no situation in which it is OK to install a mount like this. Steve had installed this flaky and shaky mount to optimize the beam size on the OL QPD.

Everyone in the lab should know better. Putting in something like this is just like sabotage - it creates extra noise in our interferometer in a sneaky way and just makes locking harder. All mounts for anything useful (including QPDs) must have highly rigid mounts. 

Use the example from the PSL relayout: use the 3/4" steel mounts and the wide aluminum bases from Newport. No more art projects using home made mounting crap, Steve.

  7506   Mon Oct 8 21:42:17 2012 JenneUpdateIOOPZT diagnosis - not fixed yet, possible solution

Note to self:

The ENV-40 amplifiers that we have supply -10V through +150V .... so don't exceed those limits.

piezojena link

  7505   Mon Oct 8 18:45:48 2012 JenneUpdateIOOPZT diagnosis - not fixed yet, possible solution

After the fuse-blowing fiasco earlier this afternoon, Koji and I took another look at the PZT controllers.

We put an ammeter in place of the fuse, and watched the current as we turned on the transformer module.  The steady-state current with no other modules plugged in is ~15mA.  However, there is a surge current right when you turn on the box which sometimes goes as high as 330mA.  Since the fuse is 250mA, this explains the fuse blowing, even though Koji had already checked out the low voltage path.

The high voltage line was connected, with +180V to the HV out pin of the backplane connector, and the (-) terminal of the power supply connected to signal ground on the board.

We inserted the PITCH module for PZT2, and we started with ~10V as our "high" voltage, and slowly increased the value (current at this time was ~60mA).  We also had a function generator plugged into the "MOD" input, which is where the epics slider goes, so that we should see a changing output voltage.  We never saw a changing output voltage.  Increasing the HV power supply didn't help. 

When Koji spun the "DC offset" knob really fast and then stopped, sometimes the output voltage as measured on the connector-converter board between the white and red wires would jump up, and then settle back down. It came back to the same value that it always was, but it was bizzarre that it would jump like that.  We suspect that that knob is an offset for use with the closed loop setting, so it isn't relevant for us anyway.  Watching the MON output, the value never changed, even when Koji did his fancy knob twirling.

We switched to the other PITCH module, and watched the output voltage on the MON output.  This time, with the function generator unplugged, so no modulation input (so we were expecting a steady DC output voltage) the number on the LCD and the MON output fluctuated wildly.  We plugged in the function generator, and the fluctuations did not change in approximate amplitude or DC offset.  They kind of looked the same. 

So, we have concluded that (a) the PZT drivers don't work, and (b) we don't understand why.  Therefore, we don't know how to fix them.

With that in mind, we are thinking of totally circumventing the PZT drivers. 

I plugged in the PZT1 connector converter board, which has Koji's circuit that he made last time when PZT1 died.  I plugged the ribbon cable which goes to the PZT, and the +\- 30V power supply, and the PZT responded!  Just plugging in the power supply puts the PZTs near the center of their nominal range.  I then put a function generator on the epics inputs for pitch and yaw (one at a time), and saw the spot move around at the ~1Hz that I was applying.  Yay!

What I think I'll do for tonight - modify the other connector converter board so that I can just use 2 HV power supplies (current limited) to steer the PZT.  I set up a TV monitor next to the PZT electronics (1Y3? 1Y4?  I forget), and it's connected to output 20 of the video switch, so I can watch the AS camera and move the PZTs by hand.  Then maybe I can try to align some stuff. (Evan is coming to work tonight, so if I electrocute myself, someone will be here to call 5000)  Koji suggested buying 2 single-channel thorlabs piezo drivers, like we have on the PSL table for the FSS loop.  These take in 0-10V and output either 0-75V, 0-100V or 0-150V (depending on which setting you choose).  These cost $712 each. This would be a more permanent solution than me just sitting out there, since we could once again control PZT2 via epics.

  7504   Mon Oct 8 14:19:17 2012 JenneUpdateIOOPZT diagnosis - not fixed yet

Quote:

What I will do tomorrow (when there is someone here to rescue me if I crispy-fry myself) is solder a wire to the now open pin of the backplane connector on the HV driver board, so that we can supply an external 180V to the pitch / yaw modules (although, obviously we won't be using the burnt yaw modules as-is).  Tomorrow I'll start by applying a nice small voltage, check that things still look okay, no shorts, and then I'll slowly increase the voltage until I get to the nominal 180V.

 I connected a thick wire to pin 22 of the backplane connector of the transformer / power supply module of the PZT box.  This is the pin that +180V is supposed to go on, to be distributed to the other boards in the crate.  Last week I had drilled a hole in the front panel so the wire can come out (since no one on campus seems to have HV panel mount connectors in stock). 

While the transformer module was isolated, not touching anything else, I applied (slowly ramping up) 180V DC, and it all looked good.

When I plugged the module back into the crate (first turning off and disconnecting the HV), I blew the 250mA fuse again.  No HV yet, just the low voltage stuff that Koji had fixed last week.  :( 

We're now out of 250mA fuses, we're supposed to get a box of them tomorrow.

  7503   Mon Oct 8 12:34:52 2012 DenUpdateModern Controlstate estimation

Quote:

 

 I guess that the estimated state has the same low pass filter, effectively, that we use to low pass the feedback signal in SUSPOS. I wonder if there is an advantage to the state estimation or not. Doesn't the algorithm also need to know about the expected seismic noise transmission from the ground to the optic?

 I think state estimation and optimal control are two different techniques that are often used together. Sometimes (as for pendulum) we can use LQG without state estimation as we need only position and velocity. But for more complex systems (like quad suspension) the states of all 4 masses can be reconstructed in some optimal way using information from only one of them if the dynamics is sufficiently well known. When current system states are measured/estimated we can apply control where all our filters are hidden.

 The algorithm needs to know about expected seismic noise transmission from the ground to the optic, but it might be not very precise. I gave it some rough estimate, there are better ways to do it. I think that we'll understand whether we need state estimation or not when we'll move to more complex systems. Brett uses a similar approach for his modal control. Interesting if these methods + seismometer readings will be able to say if one of your sensors is noisier then others.

 

 

 

 

 

 

 

 

 

  7502   Mon Oct 8 11:44:21 2012 JenneUpdateSUSETMX slow machine fine

Quote:

I think the ETMX slow machine might be dead.  All of the regular FE readbacks are fine, and the c1iscex FE computer looks fine, but the slow readbacks are all whited out. 

I turned off the damping loops for ETMX, since I don't have access to the watchdog disable/enable switch.  I guess checking this out will be task #1 for Monday morning.

 For lack of a better idea, I keyed the crate.  The computer came back up just fine, ETMX is happily damped again.

  7501   Mon Oct 8 11:15:53 2012 JenneUpdatePSLPMC and FSS had a weird weekend

Something bizarre-o was going on with the PMC and FSS over the weekend.  On the striptool, PMC's PZT looks like it was doing a sawtooth pattern for several hours.  I opened up the FSS screen, and the FSS SLOWDC had walked itself up to +10.  It's not supposed to get that far from 0. 

Here are some trends, so you can see what was going on.

10 day trend:  This weird behavior began ~Friday evening (FSS_SLOWDC ramps quickly).

PMC_sawtooth_FSSweird_8Oct2012_10dayTrend.png

1 day trend:  You can see the sawtooth pattern in PMC_PZT more clearly here.  It's at the same time as the FSS_SLOWDC is ramping rapidly, and the FSS_FAST is saturated.

PMC_sawtooth_FSSweird_8Oct2012_1dayTrend.png

Attachment 2: PMC_sawtooth_FSSweird_8Oct2012_1dayTrend.png
PMC_sawtooth_FSSweird_8Oct2012_1dayTrend.png
  7500   Mon Oct 8 10:56:59 2012 JenneUpdateSUSETMX slow machine dead??

I think the ETMX slow machine might be dead.  All of the regular FE readbacks are fine, and the c1iscex FE computer looks fine, but the slow readbacks are all whited out. 

I turned off the damping loops for ETMX, since I don't have access to the watchdog disable/enable switch.  I guess checking this out will be task #1 for Monday morning.

  7499   Mon Oct 8 09:51:30 2012 ranaUpdateModern Controlstate estimation

 

 I guess that the estimated state has the same low pass filter, effectively, that we use to low pass the feedback signal in SUSPOS. I wonder if there is an advantage to the state estimation or not. Doesn't the algorithm also need to know about the expected seismic noise transmission from the ground to the optic?

  7498   Mon Oct 8 09:45:28 2012 jamieUpdateComputersRebooted cymac0

Quote:

I rebooted cymac0 a couple of times. When I first got here it was just frozen. I rebooted it and then ran a model (x1ios). The machine froze the second time I ran ./killx1ios. I've rebooted it again.

For context, there's a is stand-alone cymac test system running at the 40m.  It's not hooked up to anything, except for just being on the martian network (it's not currently mounting any 40m CDS filesystems, for instance).  The machine is temporarily between the 1Y4 and 1Y5 racks.

  7497   Sun Oct 7 23:39:10 2012 DenUpdateModern Controlstate estimation

I've applied online state estimation technique using Kalman filter to LQG controller. It helps to estimate states that we do not measure. I've considered MC2 local damping, we measure position and want to estimate velocity that we need for control. We can either differentiate the signal or apply state estimation to avoid huge noise injection at high frequencies. In state estimation we need to know noise covariance, I've assumed that LID sensor noise is 0.1 nm. Though covariance can be calculated better.

In the time-domain figure C1:SUS-MC2_SUSPOS_IN1 = MC2 postion, C1:SUS-MC2_SUSPOS_OUT = MC2 velocity obtained by differentiation, 2 other channels are estimations of position and velocity.

Attachment 1: est_time.png
est_time.png
Attachment 2: est_freq.pdf
est_freq.pdf
  7496   Sun Oct 7 15:05:42 2012 DenUpdateSUS MC2 sus damping restored

Quote:

This is the third morning in a row that the MC2 was tripped.

 MC2 was tripped again. I think the answer is that watchdog's critical value was too small C1:SUS-MC2_PD_MAX_VAR = 10, so seismic could trip MC2. I've changed the value to 100.

mc2.png

  7495   Sun Oct 7 12:11:00 2012 AidanUpdateComputersRebooted cymac0

I rebooted cymac0 a couple of times. When I first got here it was just frozen. I rebooted it and then ran a model (x1ios). The machine froze the second time I ran ./killx1ios. I've rebooted it again.

  7494   Fri Oct 5 18:08:17 2012 ManasaConfigurationPSLAOM installation

Quote:

Do more investigation to understand what is causing the power reduction.

Is the alignment inadequate? Check the in-lock ccd image.

Is the incident power reduced? (by what?) Use dataviewer.

Is the AOM doing something? Is it active? Then how much power is it eating?

BY THE WAY, how the deflected beam is dumped?
If you don't have anything for blocking the 1st order beam, you have to expect Steve coming to you.

The PMC has been aligned and is all happy happy 

I have installed an  iris to dump the higher order beams deflected by the AOM. After installing the iris, I found that the PMC trans dropped to 0.58V and the PMC misaligned in pitch. So I've touched the 2 steering mirrors before the PMC. Now it is satisfactorily locked with PMC trans at 0.84.

I have also checked the alignment with AOM switched on. PMC trans drops to 0.15 with AOM on and comes back to 0.84 when AOM is switched off without losing lock .

  7493   Fri Oct 5 16:21:48 2012 SteveUpdateGeneralfreshmen visiters

40 plus freshmen visited the 40m today

Attachment 1: IMG_1693.JPG
IMG_1693.JPG
Attachment 2: IMG_1694.JPG
IMG_1694.JPG
  7492   Fri Oct 5 14:53:29 2012 JenneUpdatePSLTTFSS board not fully seated!

[Den, Jenne]

Den noticed that the -15V LED on the TTFSS board was not illuminated.  A further symptom of the FSS being funny was that the PC RMS Drive was constantly high (3.6 ish) and the FAST Monitor was very high, often saturated. 

We took the TTFSS board out, and put an extender card in, and it looked like all of the correct power is being supplied to the board (the +\- 24V LEDs on the extender card were illuminated).  Just to check, we put the board back in, and this time both +\- 15V LEDs came on.  So it looks like the board is fine, it just wasn't seated in there all the way.

Now the readbacks on the FSS screen look good (PC RMS Drive is less than 1, FAST Mon is 5ish), the MC is locked, and I think we're back in business. 

 

  7491   Fri Oct 5 14:40:55 2012 SteveUpdateVACRGA scan after power outage

Quote:

All the front end machines are back up after the outage.  It looks like none of the front end machines came back up once power was restored, and they all needed to be powered manually.  One of the things I want to do in the next CDS upgrade is put all the front end computers in one rack, so we can control their power remotely.

c1sus was the only one that had a little trouble.  It's timing was for some reason not syncing with the frame builder.  Unclear why, but after restarting the models a couple of times things came back.

There's still a little red, but it mostly has to do with the fact that c1oaf is busted and not running (it actually crashes the machine when I tried to start it, so this needs to be fixed!).

 The RGA was not effected by the short power outage.

Attachment 1: afterpoweroutaged10.png
afterpoweroutaged10.png
Attachment 2: pd73m10d.png
pd73m10d.png
  7490   Fri Oct 5 11:11:00 2012 DenUpdatePEMreadout box power

Quote:

Guralp readout box received +13.7 /0/ -15V instead +15V because of the broken fuse. Power provided by the source is normal.

Edit by Den: I've found a similar fuse on one of the tables and borrowed it. Guralp is not working again.

 I've meant Guralp is NOW working again 

  7489   Fri Oct 5 04:34:31 2012 ranaUpdateSUSHow about the slow machines?

Quote:

Quote:


Based on the elog entries, I assume they have not been burtrestored...

 Do you know how to burtrestore or restart slow machines?

Edit by Den: I did burtrestore of c1psl.snap from 2 days ago. Still slow machines behave not normal. For example, if I sweep C1:PSL-FSS_SLOWDC, SLOW monitor value does not change.

 Problems with Slow Machines?

Read ELOG

  7488   Fri Oct 5 01:36:49 2012 DenConfigurationPEMchanged PEM DQ channels

Quote:

We should do this wherever possible so that our channels will have real calibrations associated with them.

Next we should up the rate at which the model runs up to 16 kHz so that we can record the microphones at 16 kHz. FM radio has information up to 20 kHz. AM radio goes up to ~8 kHz. We should be at least as modern as AM radio. How do we make the change? How do we make sure the FOTON file stays OK?

 I've added calibration gains to Guralp (to um/sec) and EM172 (to Pa) channels.

We can run PEM at 16 kHz. I think Foton file stores both sos-representation and filter commands which are independent of the sampling frequency, so it should be possible to change model sampling frequency quickly.

In fact, we can save data at 64 kHz from iop models. I've done this once with MC_F channel. However, I did not test EM172 noise at frequencies > 1 kHz.

  7487   Fri Oct 5 00:29:34 2012 DenUpdatePEMreadout box power

Guralp readout box received +13.7 /0/ -15V instead +15V because of the broken fuse. Power provided by the source is normal.

Edit by Den: I've found a similar fuse on one of the tables and borrowed it. Guralp is not working again.

  7486   Thu Oct 4 23:01:49 2012 RijuparnaConfiguration cavitymode scan

 Here I am attaching the schematic diagram of the experimental set-up for IMC cavitymode scanning. A 30- 45MHz scanning signal generated by Agilent 4395A network analyzer enters EOM, which in turn modulates the laser beam entering IMC. The cavity response can be verified from reflected/transmitted beam.

I worked with the reflected beam last days. But I got no clue about the percentage of  reflected light reaching the photodiode and also the photodiode response. I would like to measure the power reaching photodiode and also would like to perform the test with transmitted beam - on wednesday if possible.

 

Attachment 1: diagram1.pdf
diagram1.pdf
  7485   Thu Oct 4 22:35:16 2012 DenUpdateSUSHow about the slow machines?

Quote:


Based on the elog entries, I assume they have not been burtrestored...

 Do you know how to burtrestore or restart slow machines?

Edit by Den: I did burtrestore of c1psl.snap from 2 days ago. Still slow machines behave not normal. For example, if I sweep C1:PSL-FSS_SLOWDC, SLOW monitor value does not change.

  7484   Thu Oct 4 22:27:54 2012 KojiUpdateSUSHow about the slow machines?

One terrible concern of mine is that the slow machines were rebooted at the power interruption.
Based on the elog entries, I assume they have not been burtrestored...

If this is true, they may cause some weird behaviors of the PSL/IOO electronics.

 

  7483   Thu Oct 4 22:16:40 2012 DenUpdatePSLPMC is locked

PZT monitor is not lying to us, I've measured it with a voltmeter. But PMC SERVO is still interesting. If I break the loop after PMCERR signal monitor (C1:PSL-PMC_BLANK=0), I can change PZT voltage from 0V (C1:PSL-PMC_RAMP = -7.3) to 132V (C1:PSL-PMC_RAMP=-10.0). If I break the loop by enabling TEST1, PZT voltage goes up to 294 V, though voltage on the TEST1 and MIXER MON is 0.

  7482   Thu Oct 4 22:16:28 2012 KojiConfigurationPSLAOM installation

Do more investigation to understand what is causing the power reduction.

Is the alignment inadequate? Check the in-lock ccd image.

Is the incident power reduced? (by what?) Use dataviewer.

Is the AOM doing something? Is it active? Then how much power is it eating?

BY THE WAY, how the deflected beam is dumped?
If you don't have anything for blocking the 1st order beam, you have to expect Steve coming to you.

  7481   Thu Oct 4 20:57:43 2012 ManasaConfigurationPSLAOM installation

Quote:

Quote:

Jenne and Jamie also find that the PMC is behaving very weird 

 Can someone detail what "weird" means? Is it singing old songs from Guns & Roses?

 It isn't singing Jan..it's dancing between 0.7 to 0 and we are not able to figure out whose the DJ ; there seems to be something else that is controlling the PMC as there is no coordination between what we do (tweaking the mirrors) and what we observe (the PD signals).

  7480   Thu Oct 4 18:48:04 2012 janoschConfigurationPSLAOM installation

Quote:

Jenne and Jamie also find that the PMC is behaving very weird 

 Can someone detail what "weird" means? Is it singing old songs from Guns & Roses?

  7479   Thu Oct 4 17:54:59 2012 ManasaConfigurationPSLAOM installation

Quote:

After the AOM work the beam wasn't well aligned to the PMC. The PMC REFL CCD shows large misalignment in yaw.

 {Jan, Manasa, Den}

We wanted to align the PMC and followed Koji's procedure detailed to us by mail. We touched the 2 steering mirrors in front of the PMC for alignment.

- Stand in front of the PMC.
- Find an oscillosocpe on the shelf in the PSL enclosure.
- This has two signals connected. One is the PMC refl dc.
  The other is the PMC trans dc.
- Minimize the refl. Maximize the trans.
- You have the CRT monitor on the MC chamber.
- Project the image of the PMC refl CCD.
  This should show some what symmetric image like an LG mode.
- Use the dataviewer to see how C1:PSL-PMC_PMCTRANSPD is recovered.

We were able to obtain 0.7 at PMC trans; but the PMC was never really stable dropped from 0.7 to 0 abruptly from time to time.

Jenne and Jamie also find that the PMC is behaving very weird 

Summary: Problem unresolved 

 

  7478   Thu Oct 4 14:08:49 2012 jamieUpdateSUSsuspensions damped

All suspension damping has been restored.

  7477   Thu Oct 4 14:04:21 2012 jamieUpdateCDSfront ends back up

All the front end machines are back up after the outage.  It looks like none of the front end machines came back up once power was restored, and they all needed to be powered manually.  One of the things I want to do in the next CDS upgrade is put all the front end computers in one rack, so we can control their power remotely.

c1sus was the only one that had a little trouble.  It's timing was for some reason not syncing with the frame builder.  Unclear why, but after restarting the models a couple of times things came back.

There's still a little red, but it mostly has to do with the fact that c1oaf is busted and not running (it actually crashes the machine when I tried to start it, so this needs to be fixed!).

  7476   Thu Oct 4 08:39:58 2012 SteveUpdateGeneralpower outage

There had to be a power outage. Laser and air condition turned back on. The vacuum is OK

Sorensen DC power supplies were tripped, so they were reset: at AUX OMC South 18V and 28V for RF PS and at 1X1 24V

 

Power Outage confirmed:

** Notification **

 

CALIFORNIA INSTITUTE OF TECHNOLOGY

                 FACILITIES MANAGEMENT

 

**PLEASE POST**

 

 

Building:         Campus

 

Date:             Thursday October 04,2012

 

This morning at 2:17 a.m. much of the City of Pasadena including our Campus experienced a electric power sag of short duration, approximately 1/10 of a second. The cause was a fault on one of Pasadena’s 17KV circuits. Some sensitive equipment have been impacted.

                 

Contact:          Mike Anchondo x-4999

 

Attachment 1: Oct4R2012.png
Oct4R2012.png
  7475   Thu Oct 4 01:06:52 2012 JenneUpdateIOOPZT diagnosis

[Koji, Jenne]

We naively hoped that just replacing the fuses would fix the problem with the PZT HV drivers.  Alas, this was not the case. 

All of our investigations (other than visual inspections) today have been of the PZT2 module.  We have not applied any electricity to any PZT1 components/modules today.

After blowing a few more fuses (not good, we know, but we really didn't know what was going on at the time and were convinced that our changes between fuse installations should prevent fuse-blowing, including removing all modules except the HV driver), we found that the YAW driver for both PZT1 and PZT2 has severe discoloration on the PCB, and several resistors and other solder joints are damaged near some high voltage regulators. Pitch on PZT1 looks a tiny bit discolored, but doesn't look totally cooked like the 2 YAW modules do.  So, at least PZT1's Yaw was cooked before we started replacing fuses, since we haven't plugged it in yet today.

We then began some more methodical checks:

We bypassed the fuses by applying 10 Vpp = ~7.2 Vrms to the input side of the big transformer on the PZT2 HV driver board.  (This usually sees the 120 Vrms from the wall AC, so we were looking at things with a factor ~16 attenuation from what they normally see.)  We then measured things on the other side of the transformer, and made sure that they made some sense (one path for 5V stuff, one path for 15V stuff, one path for 180V stuff).  One of the rectifying diode bridges (the one for HV) didn't seem to be working, and didn't seem to have all of its pins connected, as if perhaps one or more diodes inside was destroyed.

When I went home for dinner, Koji continued looking at the low voltage supply capability of the PZT2 driver.  He removed the diode bridge from the HV path, and also removed the FET that lives on the output side of the HV driver board.  He was then able to energize the HV driver and the non-burnt pitch module.  So the +\-5 V and +\-15 V paths have been confirmed okay for PZT2's driver stuff.

What I will do tomorrow (when there is someone here to rescue me if I crispy-fry myself) is solder a wire to the now open pin of the backplane connector on the HV driver board, so that we can supply an external 180V to the pitch / yaw modules (although, obviously we won't be using the burnt yaw modules as-is).  Tomorrow I'll start by applying a nice small voltage, check that things still look okay, no shorts, and then I'll slowly increase the voltage until I get to the nominal 180V.

Since the low voltage stuff on the driver board is working, once we supply an external 180V (if successful), we should be able to re-install the PZT driver and drive PZT2. 

Since both Yaw modules that we have are burnt, I am proposing that we use the PZT2 HV board (which has been checked and modified this evening) with the 2 pitch modules.  Since we are not actively utilizing the strain gauge sensors, the fact that the calibrations on these modules are not exactly the same (rather, that PZT1's pitch is not the same as PZT2's yaw) should not matter at all.  This means that we will not be able to energize PZT1 at all, but that shouldn't be a problem.  Even when PZT 2 was working, PZT1 had very, very, very limited motion through the full range of applied voltage, so having no driver connected shouldn't have an impact.

 

  7474   Wed Oct 3 23:36:54 2012 KojiConfigurationPSLAOM installation

After the AOM work the beam wasn't well aligned to the PMC. The PMC REFL CCD shows large misalignment in yaw.

Attachment 1: PMCTRANS.png
PMCTRANS.png
  7473   Wed Oct 3 23:24:05 2012 KojiUpdateSUS MC2 sus damping restored

Quote:

This is the third morning in a row that the MC2 was tripped.  Would you look at it Koji?

This may be cause by the impact of crazy WFS signal after the lock loss.
The auto locker is not fast enough to shut the WFS down before the mirrors are kicked.

Jenne and I discussed the issue and agreed that this can be solved by implementing
the same triggering algorithm as the LSC triggers.

Give us a bit more time to work on this.

  7472   Wed Oct 3 18:45:51 2012 jamieUpdateIOOwiki page for active IO tip-tilts

I made a wiki page for the active IO tip-tilts.  I should have made this a long time ago.

  7471   Wed Oct 3 16:52:16 2012 ManasaConfigurationPSLAOM installation

{Jan, Manasa}

We set start to check the performance of the AOM on the PSL table. The AOM driver spits out ~1.5W rf at 80MHz for 1V DC at its modulation input. In order to align the AOM, we reduced the input power to the AOM to ~10% using the QWP between the PBS and the laser. We touched the steering mirror before the AOM...but did not succeed in getting any appreciable first order deflection. We then released the AOM mount and moved it a few microns in and out until we obtained a significant change in power along the zero-order beam from 400mV to 100mV when the rf power was changed from 0 to ~1.5W (by changing modulation input from 0 to 1V).  The AOM was clamped at this alignment and the QWP was rotated to give maximum input power. 

During the course of aligning the AOM, the PMC unlocked and was restored after the alignment. 

All went well without having to make any emergency calls to anyone

We will now have to think about switching the AOM on and off for ringdown measurements. This could be done by either using a high-power rf switch or by switching the modulation DC input between 0 and 1V; whichever will be more comfortable to take many many ringdown measurements.

 

  7470   Wed Oct 3 16:26:58 2012 ManasaUpdatePEMants on the PSL table

I spotted around 4 within 30 minutes working at the PSL table even after the deathly spray. They seem to be running down from the cables on the oscilloscope rack to the table and the optics.

  7469   Wed Oct 3 15:58:57 2012 SteveConfigurationIOOSOS coil drivers moved

The SOS coil drivers (Atm2) were moved from 1X1 to 1Y2 location. Is this the best place to locate the  IOO Tip-Tilt steering that will replace the PJ-PZT ?

See 40m wiki T-T

Attachment 1: SOScoildriversTT.jpg
SOScoildriversTT.jpg
Attachment 2: IMG_1680.JPG
IMG_1680.JPG
  7468   Wed Oct 3 15:37:04 2012 SteveUpdatePEMants on the PSL table

We observed one or two ants climbing over PMC optics without booties and safety glasses.

The floor was mopped with strong Bayer Home Pest Control solution in the Vertex area.

Do not work inside the 40m lab if you are sensitive to chemicals!

 

Attachment 1: IMG_1690.JPG
IMG_1690.JPG
ELOG V3.1.3-